OpenStudy (anonymous):

Which of the following ratios is not equivalent to 6:10? 3/5 9/15 48/80 2445

OpenStudy (anonymous):

24/45

OpenStudy (anonymous):

I'd say the last one, 24/45 but I could be wrong

OpenStudy (anonymous):

Sht i wouldnt care aslong as i get this exam done

OpenStudy (anonymous):

Lmao, online courses?

OpenStudy (anonymous):

Yea so wack

OpenStudy (anonymous):

When we convert all in their simplest fractions 6:10-6/10=3/5 9/15=3/5 (9and 15 have common factor of three.) 48/80 have common factor of 16. Hence 3/5. 24/45 have common factor 3. Simplifying it will return 8/15. Hence, 24/45 is not an equivalent of 3/5

OpenStudy (anonymous):

Lol. At least you have access to anything on the internet. So a multitude of resources.. I find that way better than one textbook and an instructor :b Instead of just depending on whatever sources you're given, you can go and seek out what you need yourself.

OpenStudy (anonymous):

Yayyyy I was correct :D

OpenStudy (anonymous):

Thanks @Klmusta , @AngelCriner True, This is the only website my school hasnt blocked yet so im taking this to the fulest lol

OpenStudy (anonymous):

I will be greatly depressed if that happens o; I didn't know they could do that >.< lol What school do you attend?